Chapter 9 - Economic Growth II – Solow with technology

What changes when we add in technology?

(Show graph)
Answer #1

When we add technology in solow model , then breakeven investment increases.

Earlier , break even investment = (d+n)k where ,d is depreciation rate and n is population growth rate.

When there is technological progress, Break even investment = (g+d+n)k where ,g is technological growth rate.

Without technological progress, Output per worker grows at rate 0. But by addition of technology , output per worker grows at rate of g.

Earlier, total output grows at rate of n only . Now, by additon of technology total output grows at rate of (n+g).

Earlier,At steady state , sy = (d+n)k

With addition of technology , At steady state : sy = (g+d+n)k

Because of technological progress , capital per effective worker decreases.

It is shown in the following figure:

k-wen メg apital